100ml ,1 M OF sucrose solution when treated with charcol becomes 0.8M .Amount of sucrose adsorbed in moles is ?


1
Expert's answer
2020-09-16T05:12:00-0400

100 mL = 0.1 L

Initial amount in moles:

1mol/L * 0.1L = 0.1 mol

Final amount in moles:

0.8mol/L * 0.1L = 0.08 mol

Therefore, the amount of sucrose adsorbed is 0.1 - 0.08 = 0.02 mol

Answer: 0.02 moles